# Efficient Reinforcement Finetuning via Adaptive Curriculum Learning

### v2

@@ A.1 Training Configuration
A.2 AdaRFT Parameters
To enable effective AdaRFT, we incorporated a simple curriculum learning mechanism that dynamically adjusts task difficulty during training. We tuned several key hyperparameters: training batch size , target reward , sensitivity parameter , step size , and initial difficulty . These settings were used uniformly across all experiments listed in Table 3. The difficulty is updated as a function of the discrepancy between the running average reward and the target reward , using the following rule:
This formulation promotes stable learning by providing approximately linear updates when , and saturated updates when rewards deviate significantly, thereby avoiding abrupt difficulty shifts. Because the reward is bounded in and the difficulty metric spans , we set the step size to align their scales. The modulation parameter ensures smooth and controlled progression throughout training.
@@ A.2 AdaRFT Parameters
A.3 Prompt for Difficulty Estimation Using LLM as a Judge
The prompt used for difficulty estimation (as described in Section 4.1) is shown in Table 4, Table 5, and Table 6. The descriptions of the difficulty scales and examples are sourced from the AoPS Wiki.555https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings Although GPT-4o is prompted to rate problem difficulty on a scale from 1 to 10, we found that over 95% of the problems fall within the range of 1 to 5. Therefore, we clip the scores and use a revised scale from 1 to 5. In addition to integer scores, we also allow half-point increments such as 1.5, 2.0, and 2.5 for finer-grained difficulty estimation.

### v3

@@ B.1 Training Configuration
B.2 Deriving the Target-Difficulty Update Rule from a Linear Mapping
A central component of our curriculum mechanism is the update of the target difficulty based on the model’s observed reward performance. While the final update rule (Eq. 13) involves a hyperbolic tangent, it is in fact a smooth and stabilized variant of a standard linear mapping between reward space and difficulty space. We derive it here for clarity.
Step 1: Linear mapping between two intervals.
The classical linear interpolation formula for mapping a value to a target interval is
If we directly map the average reward to the difficulty range , we obtain
In our main setting, , , and , so the naive mapping simplifies to
Step 2: Mapping reward deviation instead of absolute reward.
For curriculum learning, we do not wish to reassign a new difficulty level at every step. Instead, we aim to adjust the current target difficulty depending on whether the model is performing above or below a desired target success rate . We therefore consider the deviation
Given , the deviation satisfies With the common choice , this becomes
Applying the linear mapping rule (5) from the deviation range to a symmetric difficulty-change interval yields
Thus the naive linear controller becomes
This already captures the desired behavior: difficulty increases when performance exceeds the target, decreases when performance falls short, and remains stable when .
Step 3: Stabilizing the update via a smooth saturating nonlinearity.
A purely linear controller may cause excessively large changes when the reward deviation is large or noisy. To obtain a stable update rule, we replace the linear term with a smooth, odd, saturating nonlinearity. The hyperbolic tangent is a natural choice: it behaves linearly near zero (which recovers the linear mapping) and saturates as its argument grows.
We therefore define a smoothed difficulty adjustment
Here, sets the maximum update magnitude and controls the sensitivity around the target reward. For small deviations, , so locally
recovering a linear controller with effective slope while ensuring global boundedness.
Step 4: Clipping to the valid difficulty range.
To ensure the target difficulty remains within the observed range of the data, we apply a final clipping:
The full update rule (13) is therefore a direct, smoothed generalization of a naive linear mapping between reward deviations and difficulty adjustments. It preserves the intuitive behavior of the linear controller near the target reward, while the saturating nonlinearity and clipping ensure stable, bounded, and data-consistent curriculum updates.
Because the reward is bounded in and the difficulty metric spans , we set the step size to align their scales. The modulation parameter ensures smooth and controlled progression throughout training.
@@ B.2 Deriving the Target-Difficulty Update Rule from a Linear Mapping
B.3 Prompt for Difficulty Estimation Using LLM as a Judge
The prompt used for difficulty estimation (as described in Section 4.1) is shown in Table 4, Table 5, and Table 6. The descriptions of the difficulty scales and examples are sourced from the AoPS Wiki.111https://artofproblemsolving.com/wiki/index.php/AoPS_Wiki:Competition_ratings Although GPT-4o is prompted to rate problem difficulty on a scale from 1 to 10, we found that over 95% of the problems fall within the range of 1 to 5. Therefore, we clip the scores and use a revised scale from 1 to 5. In addition to integer scores, we also allow half-point increments such as 1.5, 2.0, and 2.5 for finer-grained difficulty estimation.
Appendix C The Use of Large Language Models for ARR Jan 2026
In this ARR submission, large language models (LLMs) were used solely as writing aids for grammar correction, wording refinement, and text polishing. They were not employed for idea generation, technical contributions, or any aspect of the research beyond enhancing readability and clarity.